# UniSage: A Unified and Post-Analysis-Aware Sampling for Microservices

### v1

@@ Abstract.
Traces and logs are essential for observability and fault diagnosis in modern distributed systems. However, their ever-growing volume introduces substantial storage overhead and complicates troubleshooting. Existing approaches typically adopt a sample-before-analysis paradigm: even when guided by data heuristics, they inevitably discard failure-related information and hinder transparency in diagnosing system behavior. To address this, we introduce UniSage, the first unified framework to sample both traces and logs using a post-analysis-aware paradigm. Instead of discarding data upfront, UniSage first performs lightweight and multi-modal anomaly detection and root cause analysis (RCA) on the complete data stream. This process yields fine-grained, service-level diagnostic insights that guide a dual-pillar sampling strategy for handling both normal and anomalous scenarios: an analysis-guided sampler prioritizes data implicated by RCA, while an edge-case-based sampler ensures rare but critical behaviors are captured. Together, these pillars ensure comprehensive coverage of critical signals without excessive redundancy. Extensive experiments demonstrate that UniSage significantly outperforms state-of-the-art baselines. At a 2.5% sampling rate, it captures 56.5% of critical traces and 96.25% of relevant logs, while improving the accuracy (AC@1) of downstream root cause analysis by 42.45%. Furthermore, its efficient pipeline processes 10 minutes of telemetry data in under 5 seconds, demonstrating its practicality for production environments.
@@ 1. Introduction
Over the past decade, microservice architectures, which decompose monolithic applications into loosely coupled, independently deployable services, have become foundational to modern software systems (Ali21Survey, ; huang2021sieve, ; steam, ). This architectural paradigm offers significant benefits, such as enhanced scalability and development agility, and has been widely adopted by industry leaders, including Amazon, Google, and Netflix (practice2017microservices, ).
@@
However, as these systems grow in scale and complexity, understanding their behavior becomes increasingly challenging. Observability data (e.g., traces, logs, and metrics) play a vital role in addressing this challenge. Distributed traces record the end-to-end execution of individual requests across components, revealing causal relationships and latency bottlenecks (dapper, ; liu2023prism, ). Logs provide detailed, timestamped records of discrete events and are often used to understand control flow and error messages (logsurvey, ; jiang2024large, ). Metrics, as aggregated numerical indicators of system performance (such as CPU usage or request latency), are useful for detecting performance anomalies (Hades, ).
@@
Figure 1. A real example of the end-to-end execution of a user request in a microservice system and its observability data.:
@@
The sheer volume of this observability data, however, presents a significant hurdle, imposing substantial storage and processing overhead that makes retaining all data impractical (TraStrainer, ). For instance, a large-scale e-commerce system at Alibaba generates approximately 20 petabytes of traces daily (mint, ). While data compression can alleviate storage costs, it fails to address the inherent imbalance in observability data: the vast majority of recorded system behaviors are normal and repetitive, offering limited diagnostic value. Conversely, edge cases and failure indicators, though rare, are critically important. This disparity motivates sampling techniques, which selectively retain a representative and informative data subset, prioritizing anomalous or diverse behaviors, to reduce overhead while preserving analytical value (sifter, ; tracemesh, ; TraStrainer, ; steam, ; mint, ).
@@
Among these observability data types, trace sampling has been extensively studied because traces record fine-grained request execution paths, making them both highly valuable for debugging and extremely costly to store in full. In contrast, metrics are inherently low-dimensional and aggregated over time intervals, and therefore typically do not require sampling (micrObser, ). Logs are rich in semantics and widely used in failure diagnosis; they also incur substantial storage overhead (logshrink, ; logreducer, ). In practice, many logs are repetitive or carry limited diagnostic value, suggesting the need for selective retention. Despite this, systematic approaches to log sampling remain, to the best of our knowledge, have not been well explored.
@@
Current trace sampling strategies are primarily categorized as head-based or tail-based approaches. As illustrated in Figure 1, a trace begins when a user request enters the system and spans are generated as the request propagates through various services. Once all spans are collected, the trace is finalized and is either discarded or stored, depending on the sampling decision. Head-based sampling makes this decision at the start of this lifecycle. State-of-the-art frameworks like Jaeger (Jaeger, ) and Zipkin (Zipkin, ) employ this method, typically using low, random sampling rates (e.g., 0.1%) due to the unpredictability of failures, which often results in missing critical problematic traces. Conversely, tail-based sampling decides at the trace’s completion, allowing for more informed choices based on the full trace information. While numerous tail-based methods aim to reduce storage (huang2021sieve, ; sifter, ; TraStrainer, ; tracemesh, ; TracePicker, ), they predominantly follow a sample-before-analysis paradigm: data is sampled first, and only the retained subset undergoes downstream analysis (Figure 2 (left)). This sequence risks discarding crucial failure symptoms before any diagnosis, thereby limiting the effectiveness of troubleshooting.
@@
To overcome these limitations, we advocate for a post-analysis sampling strategy (Figure 2 (right)) that reverses the conventional sample-before-analysis paradigm. Instead of discarding data blindly or heuristically at the outset, this strategy first applies lightweight anomaly detection on the complete data stream and, when anomalies are indicated, performs root cause analysis (RCA). Only after this preliminary diagnostic step does the system selectively preserve a minimal yet highly informative subset of traces and logs for long-term storage or detailed inspection. As shown in Figure 2, this design brings several advantages: it takes anomaly relevance into consideration, leverages the full observability stream to improve diagnostic accuracy, and grounds sampling decisions in explicit analysis results, thereby offering interpretability and transparency.
@@
However, despite its promise, we identify three requirements in implementing post-analysis-aware sampling in microservice systems: (1) it must remain lightweight for production use, (2) achieve sufficiently accurate guidance by leveraging diverse signals, and (3) flexibly support both anomalous and normal scenarios. To this end, we introduce UniSage, a unified and post-analysis-aware sampling framework for microservice systems. For the efficiency and scalability requirement, we design a lightweight analysis pipeline: RCA is triggered only when anomalies are detected, and the RCA itself is based on statistical methods rather than model training. This makes the framework easy to use, model-agnostic, and decoupled from the sampling component; any RCA approach that outputs service-level anomaly scores can be seamlessly substituted. For the diagnostic accuracy requirement, UniSage integrates signals from three modalities (i.e., traces, logs, and metrics) to improve analysis robustness and broaden fault coverage. Moreover, even in the presence of false positives, the impact on sampling remains limited, since only a small fraction of additional data would be retained. Finally, to handle both anomalous and normal scenarios in a unified manner, we propose a dual-pillar sampling strategy: the edge-case-based sampler acts as a safeguard by focusing on topological and behavioral deviations, while the analysis-guided sampler leverages RCA results to retain anomaly-relevant data. Together, these pillars ensure comprehensive coverage of critical signals without excessive redundancy.
To evaluate the effectiveness and efficiency of UniSage, we conduct experiments on two public datasets (Nezha, ) derived from the widely-used TrainTicket (traintickt, ) and Online Boutique (OnlineBoutique, ) benchmarks. Results show that UniSage outperforms state-of-the-art baselines across efficiency, scalability, and diagnostic accuracy. In terms of sampling quality, it achieves 56.5% trace and 96.25% log coverage at only a 2.5% sampling rate, ensuring that anomaly-related and rare data are effectively preserved. For RCA accuracy, UniSage boosts $AC@1$ by 42.45% over the best baseline on average. On efficiency and scalability, the entire pipeline processes 10 minutes of telemetry data within 5 seconds, and reduces total runtime on large-scale datasets by 36.3% compared to the best baseline, demonstrating strong practicality for production deployment.
The key contributions of this paper are as follows:
  • We propose a post-analysis-aware sampling paradigm that overcomes the limitations of traditional sample-before-analysis approaches. Instead of discarding data blindly or heuristically, it incorporates analysis results into sampling decisions to retain the most valuable data for long-term storage.
  • We design and implement UniSage, the first unified analysis and sampling framework for both traces and logs. It features an efficient anomaly analysis module and a dual-pillar sampling strategy, effectively balancing diagnostic accuracy, efficiency, and sampling coverage across anomalous and normal scenarios.
  • We conduct extensive experiments on two widely-used microservice benchmarks, demonstrating that UniSage identifies more valuable traces and logs under the same budget and improves downstream root cause analysis performance compared to state-of-the-art methods.
@@ 2. Background and Motivation
2.1. Background
2.1.1. Observability
Originally introduced in control theory (kalman1960general, ), observability refers to the extent to which a system’s internal states can be inferred from its external outputs. In modern software systems, this concept has been adapted to describe the ability to understand system behavior through telemetry data, including metrics, logs, and traces (opentelemetry, ). These signals are commonly referred to as the “three pillars” of observability (ObservabilityBook, ), collectively enabling engineers to monitor complex distributed systems (yu2024monitorassistant, ), diagnose failures (huang2024demystifying, ; jiang2025l4, ), and ensure reliable operation.
Traces records the execution process of a user request, which typically involves multiple operations across different services. Each operation is represented as a span, and a complete trace is composed of one or more spans organized as a directed acyclic graph (DAG), where edges capture the parent-child (caller-callee) relationships between spans. As shown in Figure 1, a user request flows through multiple microservices, each of which generates observability data such as traces, logs, and metrics during execution. These signals are collected and stored in a collector for subsequent analysis. The trace representation of this request is illustrated in the right part of Figure 1. Each span records metadata such as the service and operation (i.e., span name), and duration.
Metrics represent time series data points that track key performance indicators (KPIs), such as system error rate, CPU utilization, and request rate for a given service. The primary objective of metrics is to provide insights into the health and operational status of a specific component or system. Owing to their compact nature, metrics can be efficiently collected at scale, which is particularly advantageous in large, distributed environments. Furthermore, metrics are well-suited for pre-aggregation at the component level prior to collection, thereby minimizing the computational overhead associated with processing and storing (micrObser, ).
Logs are runtime messages generated by services or components to record system events, often categorized by verbosity levels such as info, warn, and error. Unlike traces, logs are not inherently linked to specific user requests or transactions, but have long served as a primary tool for developers and operators to understand system behavior. In modern observability setups, however, logs generated during the lifecycle of a request are often instrumented with the corresponding trace_idtrace\_id, enabling direct correlation between logs and traces (e.g., the error log shown at the bottom of Figure 1 is associated with the trace illustrated above). Due to the large volume of logs generated in modern systems, log compression techniques have been increasingly adopted to mitigate storage overhead (yu2024logCompression, ; logshrink, ; liu2019logzip, ).
2.1.2. Root Cause Analysis
Root Cause Analysis (RCA), also referred to as troubleshooting, is the process of identifying the underlying causes of failures in a software system, such as service outages or performance degradations. It is considered the primary task of observability data, and many efforts have been devoted to identifying root causes in microservice systems (rcasurvey, ; Eadro, ; Nezha, ; PDiagnose, ; microrca, ; microcause, ; microhecl, ). During incident response, on-call engineers must collect and analyze telemetry data to understand the sequence of events leading to a failure. Furthermore, failures can propagate across service dependencies (e.g., a payment service depending on a database) or deployment configurations (e.g., co-located containers on a shared host), often resulting in cascading effects that obscure the true origin of the problem.
@@ 2.2. Motivation
Motivation 1: Lack of a unified sampling strategy across observability data. Modern observability relies on three key telemetry signals: metrics, traces, and logs (kalman1960general, ; opentelemetry, ). A fundamental limitation in current pipelines is the siloed approach to sampling these signals. While trace sampling has been extensively studied (huang2021sieve, ; sifter, ; tracemesh, ; steam, ) and metrics are naturally optimized through aggregation (micrObser, ), logs are typically handled by independent mechanisms that focus on compression or coarse filtering rather than diagnostic context (yu2024logCompression, ; logshrink, ; liu2019logzip, ). This decoupling of sampling decisions creates a critical gap for troubleshooting: the loss of correlated context across modalities.
@@
Motivation 3: Improving interpretability of sampling decisions. Most existing sampling methods operate as black boxes, providing little transparency into why particular traces or logs are retained or discarded. Typically driven by heuristics or learned models, these approaches lack human-understandable justifications, which undermines trust, complicates troubleshooting, and restricts operators’ ability to tune or audit the process. In failure diagnosis, however, understanding the rationale behind preserved (or discarded) data can be as critical as the data itself. To address this, we advocate incorporating interpretable signals that explicitly explain each sampling decision. These signals, such as multi-modal patterns illustrated in Table 1, capture fine-grained behavioral deviations and can serve as direct evidence for diagnosis.
3. Problem Formulation
Consider a large-scale distributed system with $N$ services, where telemetry data (i.e., logs $\mathcal{L}$, metrics $\mathcal{M}$, and traces $\mathcal{T}$) are collected individually for each service. Let $\mathbf{X}$ denote the complete set of telemetry data, with $\mathbf{X}_{\mathcal{F}}$ collected during the fault-free construction phase $\mathcal{F}$ and $\mathbf{X}_{\mathcal{P}}$ collected during the fault-suffering production phase $\mathcal{P}$. The objective is twofold: (1) perform lightweight and accurate fault diagnosis over $\mathbf{X}_{\mathcal{P}}$ using full observability data, and (2) derive a sampling strategy that retains only a small yet informative subset of data for long-term storage and further analysis.

### v2

# UniSage: A Unified and Post-Analysis-Aware Sampling Framework for Microservice System

@@ Abstract.
Traces and logs serve as the backbone of observability in microservice architectures, yet their sheer volume imposes prohibitive storage and computational burdens. To reduce overhead, operators rely on sampling; however, current frameworks generally employ a sample-before-analysis strategy. This approach creates a fundamental trade-off: to save space, systems must discard data before knowing its diagnostic value, often losing critical context required for troubleshooting anomalies and latency spikes. In this paper, we propose UniSage, a unified sampling framework that addresses this trade-off by adopting a post-analysis-aware paradigm. Unlike prior works that focus solely on tracing, UniSage integrates both traces and logs, leveraging a lightweight anomaly detection and root cause analysis module to scan the full data stream before sampling decisions are made. This pre-computation enables a dual-pillar strategy: an analysis-guided sampler that retains high-value data associated with detected anomalies, and an edge-case sampler that preserves rare but critical behaviors to ensure diversity. Evaluation on three datasets confirms that UniSage achieves superior data retention. At a 2.5% sampling rate, UniSage captures 71% of critical traces and 96.25% of relevant logs, substantially exceeding the best existing methods (which achieve 42.9% and 1.95%, respectively). Moreover, evaluations on a real-world dataset demonstrate UniSage’s efficiency; it processes a 20-minute multi-modal data block in an average of 10 seconds, making it practical for production environments.
@@ 1. Introduction
Over the past decade, microservice architectures, which decompose monolithic applications into loosely coupled, independently deployable services, have become foundational to modern software systems (Luo et al., 2021; Huang et al., 2021; He et al., 2023). This architectural paradigm offers significant benefits, such as enhanced scalability and development agility, and has been widely adopted by industry leaders, including Amazon, Google, and Netflix (Pautasso et al., 2017).
@@
However, as these systems grow in scale and complexity, understanding their behavior becomes increasingly challenging. Observability data (i.e., traces, logs, and metrics) play a vital role in addressing this challenge. Distributed traces record the end-to-end execution of individual requests across components, revealing causal relationships and latency bottlenecks (Sigelman et al., 2010; Liu et al., 2023). Logs provide detailed, timestamped records of discrete events and are often used to understand control flow and error messages (He et al., 2021; Jiang et al., 2024b). Metrics, as aggregated numerical indicators of system performance (such as CPU usage or request latency), are useful for detecting performance anomalies (Lee et al., 2023b).
@@
Figure 1. A real example of the end-to-end execution of a user request in a microservice system and its observability data.
@@
The sheer volume of this observability data, however, presents a significant hurdle, imposing substantial storage and processing overhead that makes retaining all data impractical (Huang et al., 2024a). For instance, a large-scale e-commerce system at Alibaba generates approximately 20 petabytes of traces daily (Huang et al., 2025). While data compression can alleviate storage costs, it fails to address the inherent imbalance in observability data: the vast majority of recorded system behaviors are normal and repetitive, offering limited diagnostic value. Conversely, edge cases and failure indicators, though rare, are critically important. This disparity motivates sampling techniques, which selectively retain a representative and informative data subset, prioritizing anomalous or diverse behaviors, to reduce overhead while preserving analytical value (Las-Casas et al., 2019; Chen et al., 2024; Huang et al., 2024a; He et al., 2023; Huang et al., 2025).
@@
Among these observability data types, trace sampling has been extensively studied because traces record fine-grained request execution paths, making them both highly valuable for debugging and extremely costly to store in full. In contrast, metrics are inherently low-dimensional and aggregated over time intervals, and therefore typically do not require sampling (Microsoft, 2025). Logs are rich in semantics and widely used in failure diagnosis; they also incur substantial storage overhead (Li et al., 2024; Yu et al., 2023a). In practice, many logs are repetitive or carry limited diagnostic value, suggesting the need for selective retention. Despite this, log sampling has not yet been systematically explored, to the best of our knowledge.
@@
Current trace sampling strategies are primarily categorized as head-based or tail-based approaches. As illustrated in Figure 1, a trace begins when a user request enters the system, and spans are generated as the request propagates through various services. Once all spans are collected, the trace is finalized and is either discarded or stored, depending on the sampling decision. Head-based sampling makes this decision at the start of this lifecycle. State-of-the-art frameworks like Jaeger (Jaeger, 2025) and Zipkin (Zipkin, 2025) employ this method, typically using low, random sampling rates (e.g., 0.1%) due to the unpredictability of failures, which often results in missing critical problematic traces. Conversely, tail-based sampling decides at the trace’s completion, allowing for more informed choices based on the full trace information. While numerous tail-based methods aim to reduce storage (Huang et al., 2021; Las-Casas et al., 2019; Huang et al., 2024a; Chen et al., 2024; Xie et al., 2025), they predominantly follow a sample-before-analysis paradigm: data is sampled first, and only the retained subset undergoes downstream analysis (Figure 2 (left)). This sequence risks discarding crucial failure symptoms before any diagnosis, thereby limiting the effectiveness of troubleshooting.
@@
To overcome these limitations, we advocate for a post-analysis-aware sampling paradigm (Figure 2 (right)) that reverses the conventional workflow. Instead of discarding data blindly or heuristically at the outset, this strategy first applies lightweight anomaly detection on the complete data stream and, when anomalies are indicated, triggers root cause analysis (RCA). Only after this preliminary diagnostic step does the system selectively preserve a minimal yet highly informative subset of traces and logs. As shown in Figure 2, this design brings several advantages: it prioritizes anomaly relevance, leverages the full observability stream for accuracy, and grounds sampling decisions in explicit analysis results, thereby retaining valuable data.
@@
However, realizing this paradigm faces three key challenges: (1) Efficiency: the analysis must remain lightweight enough for production deployment; (2) Accuracy: the guidance must be robust by leveraging diverse signals; and (3) Coverage: the system must flexibly support both anomalous and normal scenarios. To this end, we introduce UniSage, a unified and post-analysis-aware sampling framework for microservice systems. To address efficiency, we design a lightweight analysis pipeline where RCA is triggered only upon anomaly detection and relies on statistical methods rather than heavy model training. This makes the framework model-agnostic and decoupled from the sampling component. To improve diagnostic accuracy, UniSage integrates signals from three modalities (i.e., traces, logs, and metrics), broadening fault coverage. Even in the presence of false positives, the impact remains limited as only a small fraction of additional data is retained. Finally, to unify the handling of different scenarios, we propose a dual-pillar sampling strategy: the edge-case-based sampler safeguards topological and behavioral deviations, while the analysis-guided sampler leverages RCA results to retain anomaly-relevant data. Together, these pillars ensure comprehensive coverage without excessive redundancy.
We evaluate UniSage on three widely used multi-modal datasets (Yu et al., 2023b; Google, 2024a), including two representative microservice benchmarks (i.e., Train Ticket (Zhou et al., 2018) and Online Boutique (Google, 2023)) and one real-world industry microservice system (Google, 2024b). Extensive experiments demonstrate that UniSage consistently preserving diagnostic and edge-case traces/logs while aggressively filtering out low-value noise, achieves 71% trace and 96.25% log coverage at only 2.5% sampling rate. Moreover, UniSage improves RCA accuracy $AC@1$ by 42.45% over the best baseline on average. In terms of efficiency and scalability, evaluations on a real-world dataset demonstrate the entire pipeline processes a 20-minute multi-modal data block in an average of 10 seconds, demonstrating strong practicality for production deployment.
The key contributions of this paper are threefold:
  • Paradigm Shift: We introduce a post-analysis-aware sampling paradigm that rethinks sampling in microservice observability by performing diagnosis prior to data retention, fundamentally departing from the traditional sample-before-analysis workflow.
  • Unified Framework: We propose UniSage, a unified and lightweight framework that guides trace and log sampling using multi-modal analysis results, combining analysis-guided and edge-case-based sampling to balance diagnostic relevance and coverage diversity.
  • Comprehensive Evaluation: We extensively validate the effectiveness and efficiency of the proposed paradigm on two widely used microservice benchmarks and one real-world system, demonstrating substantial improvements over state-of-the-art methods under the same sampling budget.
@@ 2. Background and Motivation
2.1. Observability
Originally introduced in control theory (Kalman, 1960), observability describes the extent to which internal states are inferred from external outputs. In modern software, this relies on telemetry data, including metrics, logs, and traces (Foundation, 2025). These “three pillars” (Sridharan, 2018) enable engineers to monitor distributed systems (Yu et al., 2024b), diagnose failures (Huang et al., 2024b; Jiang et al., 2025), and ensure reliability. Traces record a user request’s execution path, composed of spans (individual operations) organized as a directed acyclic graph representing caller-callee relationships. As illustrated in Figure 1, traces visualize the request flow across microservices, capturing metadata such as service names, operations, and duration. Metrics are time-series data tracking Key Performance Indicators (KPIs) like error rates and CPU utilization. They provide insights into system health and, due to their compact nature, allow for efficient, scalable collection and pre-aggregation with minimal overhead (Microsoft, 2025). Logs capture discrete system events with varying verbosity levels. While traditionally standalone, modern logs are often tagged with a traceID to correlate with specific requests (e.g., the error log in Figure 1). To manage high data volumes, log compression is widely adopted to mitigate storage costs (Yu et al., 2024a; Li et al., 2024; Liu et al., 2019).
@@ 2.2. Motivation
Motivation 1: Lack of a unified sampling strategy across observability data. Modern observability relies on three key telemetry signals: metrics, traces, and logs (Kalman, 1960; Foundation, 2025). A fundamental limitation in current pipelines is the siloed approach to sampling these signals. While trace sampling has been extensively studied (Huang et al., 2021; Las-Casas et al., 2019; Chen et al., 2024; He et al., 2023) and metrics are naturally optimized through aggregation (Microsoft, 2025), logs are typically handled by independent mechanisms that focus on compression or coarse filtering rather than diagnostic context (Yu et al., 2024a; Li et al., 2024; Liu et al., 2019). This decoupling of sampling decisions creates a critical gap for troubleshooting: the loss of correlated context across modalities.
@@
2.3. Problem Formulation
We consider a microservice system $\mathcal{S}=\{s_{1},\dots,s_{N}\}$ comprising $N$ services. The system generates multi-modal telemetry data, including logs $\mathcal{L}=\{l_{1},\cdots,l_{L}\}$, metrics $\mathcal{M}=\{m_{1},\cdots,m_{M}\}$, and traces $\mathcal{T}=\{t_{1},\cdots,t_{T}\}$. The global dataset is defined as $\mathbf{X}=\{(\mathbf{X}^{\mathcal{L}}_{i},\mathbf{X}^{\mathcal{K}}_{i},\mathbf{X}^{\mathcal{T}}_{i})\}_{i=1}^{N}$. This data is partitioned into two phases: the construction phase $\mathbf{X}_{\mathcal{C}}$ (historical normal data) and the Observation Phase $\mathbf{X}_{\mathcal{O}}$ (real-time streaming data).
Our framework establishes an end-to-end pipeline with three sequential objectives. First, anomaly detection predicts a binary system state $y\in\{0,1\}$, where $y=0$ indicates normal operation and $y=1$ indicates an abnormal state. Second, if $y=1$, root cause analysis (RCA) ranks services to identify the most likely faulty components, producing an ordered list $\mathcal{R}=[r_{1},\dots,r_{N}]$, where $r_{k}$ denotes the index of the $k$-th ranked service. Third, adaptive sampling determines the sampling probabilities for traces $\mathbf{P}^{\mathcal{T}}=[p_{t_{1}},\dots,p_{t_{T}}]$ and logs $\mathbf{P}^{\mathcal{L}}=[p_{l_{1}},\dots,p_{l_{L}}]$, jointly denoted as $\mathbf{P}=\{\mathbf{P}^{\mathcal{T}},\mathbf{P}^{\mathcal{L}}\}$. The decision to retain data is based on the sampling probability. Formally, the sampling decision is defined by a conditional sampling function $f(\mathbf{X},y,\mathcal{R})\rightarrow\mathbf{P},$ when $y=0$, the sampling probabilities are determined solely by an edge-case sampler that preserves informative but rare normal data. When $y=1$, the sampling function integrates signals from both the edge-case sampler and an analysis-guided sampler derived from the RCA results $\mathcal{R}$, enabling anomaly-aware adaptive sampling. Overall, the framework maps the input data stream to
$\mathbf{X}\rightarrow(y,\mathcal{R},\mathbf{P},\mathbf{X}^{\prime}),$
where $\mathbf{X}^{\prime}$ denotes the retained subset of telemetry data after adaptive sampling.